OpenStudy (anonymous):

Help? Simplify 5+x/x+5.

OpenStudy (anonymous):

\[\frac{5+x}{x+5}\]?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

So it would just stay the same?

OpenStudy (anonymous):

=1

OpenStudy (anonymous):

\[{{5+x} \over {x+5}} = {{x+5} \over {x+5}} = 1\]